OpenStudy (anonymous):

How far from the surface of the earth does the acceleration due to gravity become 4% of its value on the surface of the earth ?

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

g'=4/100 *g GM/ (R+h)^2 = 4/100 * (GM/R^2)

OpenStudy (anonymous):

Now, simplify it

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):